OpenStudy (anonymous):

Mr. Trundle used exactly 128 ceramic tiles to cover the floor of a shower. If each tile is a square with 6-inch sides, what is the area of the shower floor?

OpenStudy (anonymous):

128 * 36 sq ins

OpenStudy (anonymous):

= 128 * 36 ------- sq ft = 128 / 4 = 32 sq ft 144